OpenStudy (anonymous):

In a given experiment 102 g of PCl5 was placed in an empty 1.0 L container. Equilibrium was established with the following equation PCl5----> PCl3+Cl2 k=1.2*10^-4 How do I find the values at equilibrium?

OpenStudy (aaronq):

>Write a balanced equation >Set up an equilibrium expression >convert mass of PCl5 to moles -> plug into equilibrium expression

OpenStudy (anonymous):

I'm not given values for the PCl3 or Cl2

OpenStudy (aaronq):

thats what they want you to find, they gave you the equilibrium constant K.

OpenStudy (anonymous):

so you would set it up as x^2/ .49-x= .00012?

OpenStudy (aaronq):

yep, solve for x

OpenStudy (anonymous):

x^2+ .00058x-.0012=0?

OpenStudy (aaronq):

doesn't look right

OpenStudy (anonymous):

how do i make it into a quadratic?

OpenStudy (aaronq):

x^2/(.49-x)= .00012 x^2= .00012(.49-x) distribute, set equal to zero

OpenStudy (anonymous):

x^2+.00012x-.000058?

OpenStudy (aaronq):

yep looks better

OpenStudy (anonymous):

thanks soo much